NEXT
Make New Entry,
Make Followup Entry
User name brads
Log entry time 13:53:16 on December 16, 2008
Entry number 253623
keyword=BB Cerenkov modifications
PMTs 4, 5, 6, 7 on the beamline side of the Cerenkov are now immediately
run through a Philips (NIM) 10x amplifier mounted right on the Cerenkov
tank. The signals then run through the 10x amplifier on the front-end
rack as usual. HVs on those PMTs were dropped to offset the extra
amplification. LED runs were taken to check the single p.e. peak
position and pedestal widths. (I'll post a few snapshots later.)
Bottom line is that the resolution is definitely degraded, but not too
badly. Hopefully this will let me run those PMTs safely under the
present conditions.
The TDC signal for PMT 7 has been missing for at least a week. I tracked
it down to a bad (half of a) summing module. I moved the signals to the
unused half of the same module and it seems to be OK now.
I remeasured the discriminator threshold on the per-PMT discriminators
and it was (still) -100mV.
A copy of this log entry has been emailed to: meziani